suppose that an airplane engine will fail, when in flight, with probability 1−p independently from engine to engine; suppose that the airplane will make a successful flight if at least 50 percent of its engines remain operativ
1
Expert's answer
2020-09-07T15:42:19-0400

Here given,

engine fail probabilityP(F) = 1−p independently from engine to engine.

engine success probability P(S) = 1-(1-p)

= p

There no question specified . I think a question is like

'For what values of p is a four-engine plane preferable to a two-engine plane?'


For making four engine airplane success at least 2 engine of them should be success.


so, probability = "(4C2)p^2(1-p)^2 + (4C3)p^3(1-p)^1 + (4C1)p^4(1-p)^0\n\n\n\n= 6p^2(1-p)^2+4p^3(1-p) + 4p^4"


required Probability for 2 engine airplane success at least 1 engine should success

= "(2C1)p^1(1-p)^1 + (2C2)p^2(1-p)^0"

=2(p)(1-p) + p2

The four-engine plane is safe if:


 "6p^2(1-p)^2+4p^3(1-p) + 4p^4 \geq 2(p)(1-p) + p^2"


"6 p(1\u2212 p) + 4 p (1\u2212 p) + p \u2265 2 \u2212 p"


"3p^3-8p^2+7p-2 \geq0 \,or \, ( p-1)^2 (3p-2) \geq 0"


"3p-2 \geq 0 
\,so, \, p \geq2/3"